OpenStudy (anonymous):

Which polynomial identity will prove that 49 - 4 = 45?

OpenStudy (anonymous):

Answer choices:

OpenStudy (anonymous):

as is clear on your equation, 49 is a square given by 7x7 as is clear in your equation, 4 is a square given by 4x4 the operator given is - ( minus) you have two squares, and a difference sign a^2 - b^2 = (a-b)(a+b) 7^2 -2^2 = (7-2)(7+2) = (5)(9) = 45 it is the diffeernce of two squares

OpenStudy (anonymous):

Thank you friend :D